the water current is 2 miles per hour . a canoe can travel 6 mile downstream with the current, in the same amount of time it travel 2 miles upstream against the current. what is the canoe rate in still water?
Answer by scott8148(6628)

t = d / r ___ times are the same, upstream and downstream

6 / (c + 2) = 2 / (c - 2)

"cross" multiplying ___ 6c - 12 = 2c + 4

4c = 16 ___ c = 4